# Exact Factorization of Unitary Transformations with Spin-Adapted Generators

### v1

@@ I Introduction
Conventional unitary coupled cluster (UCC) ansätze[UCC:rev] and their adaptive extensions[QCC, iQCC, Grimsley2019-zw, Qubit-ADAPT, Burton2023] construct the variational wavefunction from a pool of anti-Hermitian generators chosen to balance expressiveness and circuit depth. In principle, the complete set of $4^{N}-1$ traceless Pauli generators acting on $N$ qubits spans the full $\mathfrak{su}(2^{N})$ Lie algebra, guaranteeing convergence to the exact ground state within a fixed orbital basis. The exponential scaling of this set, however, makes its direct use computationally infeasible. To achieve polynomial scaling, smaller operator pools that are closed under commutation have been proposed in both fermionic[Grimsley2019-zw] and qubit operator spaces[Qubit-ADAPT]. While such reduced pools can in principle generate the full algebra, many of them break key physical symmetries, including total spin ($S^{2}$), its projection ($S_{z}$), and particle number ($N_{e}$).
Maintaining these symmetries during variational optimization is desirable not only for physical consistency but also for improved convergence and reduced redundancy in the variational parameters. Several strategies have been proposed to enforce or restore symmetry, including penalty-based constraints[ryabinkin2019constr], projection methods[ProjEA2019], and modified symmetry-conserving generators[Anselmetti_2021, Burton2023]. The DISCO framework[Burton2023], for example, introduces one-electron and electron-pair rotations that preserve total spin by construction. Although effective for spin conservation, this approach restricts the operator pool to zero-seniority excitations (pair rotations), thereby reducing the overall expressivity of the ansatz.
@@
Since the number of basis elements in the relevant Lie algebras is modest (fewer than 150), the entire factorization can be carried out efficiently in the adjoint representation using low-dimensional matrix algebra. Instead of employing symbolic Baker–Campbell–Hausdorff expansions, we reduce the reparametrization problem to a small nonlinear optimization involving matrix exponentials in the adjoint representation. This numerical treatment achieves high precision while avoiding symbolic complexity.
@@ II.1 Lie algebra framework
where $A_{k}\in\mathcal{A}$ form a Lie algebra generated by the commutator closure of the $\{A_{k}\}$ operators. The dimensionality of these algebras, $m\lesssim 138$, depends on the seniority and the specific index pattern. Each separate exponential $\exp(tA_{k})$ admits a well-known gate representation, and therefore our objective is to decompose
@@
The first question is whether such a product representation exists for the Lie algebras considered. This issue is nontrivial because certain Lie algebras contain generator sets for which the exponential of a linear combination cannot be represented as a finite product of exponentials of the individual generators [nonSurj]. In our case, all generators are anti-Hermitian operators that form subalgebras of the compact $\mathfrak{su}(N)$ algebra, for which the exponential map is surjective. Consequently, any element of the Lie group generated by the left-hand side (LHS) of Eq. (1) can be expressed by an appropriate choice of parameters in the right-hand side (RHS).
@@ II.3 Adjoint representation
To determine the parameters in Eq. (1) we use the adjoint representation, which maps each algebra element to a matrix of dimension equal to that of the Lie algebra. For the algebras considered here, the dimensions (depending on seniority and index pattern) are $5$, $28$, and $138$. These sizes are substantially smaller than those arising in faithful representations built in the computational basis of the corresponding spin orbitals. For example, using the computational basis in the Fock space of the involved spin orbitals requires matrices of dimensions $64$ and $256$ for seniorities $2$ and $4$, respectively. Therefore, we adopt the adjoint representation, in which operator matrices are constructed from the commutator (structure) constants: each element $A_{i}$ corresponds to a matrix $M^{(i)}$ defined by
@@
This objective is representation-invariant within $\exp(\mathfrak{s})$ and encodes the Baker–Campbell–Hausdorff identity for the chosen semisimple Lie algebra. In practice, we employ a smooth quasi-Newton optimizer (BFGS). Because $m\leq 138$ in all cases considered, each evaluation of the norm difference is inexpensive and convergence is rapid. Code to perform this minimisation is available on GitHub[spinadapted].
@@ II.4 Implementation of unitaries
$N_{\text{Pauli}}^{\text{JW}}$ 48 48 384 768
@@ III.1 Computational details
To evaluate the effect of spin-adapted fermionic generators on the efficiency and accuracy of variational algorithms, we carried out a series of adaptive VQE calculations using different operator pools. All calculations employed the STO-3G basis set and were converged when the maximum absolute value of the gradient of the operator pool was below $10^{-5}$ a.u. The state-vector simulator implemented in SlowQuant[slowquant] was used to perform noiseless classical emulations. Energy minimizations employed the BFGS algorithm as implemented in SciPy[2020SciPy-NMeth].
The operator pool for the conventional fermionic singles and doubles (SD) adaptive fermionic ansatz consisted of $\{\hat{G}_{i_{\sigma}}^{a_{\sigma}},\,\hat{G}_{i_{\sigma}j_{\tau}}^{a_{\sigma}b_{\tau}},\,\hat{G}_{i_{\sigma}j_{\tau}}^{a_{\tau}b_{\sigma}}\},\quad i,j<a,b$, this ansatz is equal to the ADAPT ansatz[Grimsley2019-zw]. For the spin-adapted fermionic ansatz ($\mathrm{S}_{\mathrm{SA}}\mathrm{D}_{\mathrm{SA}}$), the operator pool was $\{\hat{G}^{\mathrm{SA}}_{ia},\,\hat{G}^{\mathrm{SA}}_{ijab},\,\hat{G}^{\mathrm{SA}^{\prime}}_{ijab}\},\quad i,j<a,b$. Here, $\hat{G}^{\mathrm{SA}}_{ia}=\hat{G}_{i_{\alpha}}^{a_{\alpha}}+\hat{G}_{i_{\beta}}^{a_{\beta}}$, is a fermionic spin-adapted single excitation operator. Finally, for the spin-adapted singles and pair-doubles ansatz ($\mathrm{S}_{\mathrm{SA}}\mathrm{D}_{\mathrm{pair}}$), the operator pool was $\{\hat{G}^{\mathrm{SA}}_{ia},\,\hat{G}^{\mathrm{SA}}_{iiaa}\},\quad i<a,$ where $\hat{G}^{\mathrm{SA}}_{iiaa}$ corresponds to $\hat{G}_{i_{\alpha}i_{\beta}}^{a_{\alpha}a_{\beta}}$. Each adaptive VQE ansatz was constructed iteratively by selecting the operator with the largest energy gradient magnitude at each step until the convergence criterion was reached.
@@ III.2 Adaptive VQE calculations
Figure 1: Adaptive VQE calculations for $\text{H}_{2}\text{O}$ with $\theta_{\mathrm{HOH}}=109^{\circ}$ and $R_{\mathrm{OH}}=0.977\ \mathrm{\AA }$. SD denotes the fermionic operator pool, $\text{S}_{\text{SA}}\text{D}_{\text{SA}}$ the spin-adapted fermionic operator pool, and $\text{S}_{\text{SA}}\text{D}_{\text{pair}}$ the pool of spin-adapted singles and pair doubles.
@@
The alternative pool of spin-adapted singles and pair doubles (green dashed line) also conserves total spin and consists of circuits that can be directly implemented with known gate structures. However, this ansatz frequently becomes trapped in local minima and, in this case, converged to a solution $8.8\times 10^{-3}$ Hartree above the FCI reference. A similar shortcoming of pair-doubles-based operator pools has been reported previously by Burton et al.[Burton2023]
Figure 2: Adaptive VQE calculations for $\text{Be}\text{H}_{2}$ with $\theta_{\mathrm{HBeH}}=180^{\circ}$ and $R_{\mathrm{BeH}}=1.35\ \mathrm{\AA }$. SD denotes the fermionic operator pool, $\text{S}_{\text{SA}}\text{D}_{\text{SA}}$ the spin-adapted fermionic operator pool, and $\text{S}_{\text{SA}}\text{D}_{\text{pair}}$ the pool of spin-adapted singles and pair doubles.
@@
Previous studies have attempted to address this problem by adding penalty terms to the Hamiltonian during variational optimization[Kuroiwa2021] or by applying spin-projection techniques[Tsuchimochi2020]. However, these approaches increase computational cost either by complicating the energy functional or by requiring additional ancilla qubits. Using spin-adapted operators naturally preserves the desired spin symmetry without modifying the Hamiltonian or introducing extra qubits, offering a simple and physically consistent solution.
@@ Appendix B Seniority 4, intermediate triplet
This 138-element Lie algebra originates from the commutator closure of 6 elementary generators involved in the symmetrized generator of Eq. (II.2)
@@
$A_{41}$ $=-(\hat{n}(j_{\beta})+\hat{n}(a_{\beta})-2\hat{n}(j_{\beta})\hat{n}(a_{\beta}))\hat{G}_{i_{\alpha}j_{\alpha}}^{a_{\alpha}b_{\alpha}}$
$A_{42}$ $=-(\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})-2\hat{n}(i_{\alpha})\hat{n}(b_{\alpha}))\hat{G}_{i_{\beta}j_{\beta}}^{a_{\beta}b_{\beta}}$
$A_{43}$ $=(\hat{n}(j_{\alpha})+\hat{n}(b_{\alpha})-2\hat{n}(j_{\alpha})\hat{n}(b_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{44}$ $=(\hat{n}(i_{\alpha})\hat{n}(j_{\beta})-\hat{n}(i_{\alpha})\hat{n}(b_{\beta})-\hat{n}(j_{\beta})\hat{n}(a_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{45}$ $=(\hat{n}(i_{\alpha})+\hat{n}(a_{\alpha})-2\hat{n}(i_{\alpha})\hat{n}(a_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{46}$ $=(\hat{n}(i_{\beta})\hat{n}(j_{\alpha})-\hat{n}(i_{\beta})\hat{n}(b_{\alpha})-\hat{n}(j_{\alpha})\hat{n}(a_{\beta})+\hat{n}(a_{\beta})\hat{n}(b_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{47}$ $=(\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})-2\hat{n}(i_{\alpha})\hat{n}(b_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{48}$ $=(\hat{n}(i_{\beta})\hat{n}(j_{\alpha})-\hat{n}(i_{\beta})\hat{n}(a_{\alpha})-\hat{n}(j_{\alpha})\hat{n}(b_{\beta})+\hat{n}(a_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{49}$ $=(\hat{n}(j_{\alpha})+\hat{n}(a_{\alpha})-2\hat{n}(j_{\alpha})\hat{n}(a_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{50}$ $=(\hat{n}(i_{\alpha})\hat{n}(j_{\beta})-\hat{n}(i_{\alpha})\hat{n}(a_{\beta})-\hat{n}(j_{\beta})\hat{n}(b_{\alpha})+\hat{n}(a_{\beta})\hat{n}(b_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{51}$ $=(\hat{n}(i_{\beta})+\hat{n}(a_{\beta})-2\hat{n}(i_{\beta})\hat{n}(a_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{52}$ $=(\hat{n}(j_{\beta})+\hat{n}(b_{\beta})-2\hat{n}(j_{\beta})\hat{n}(b_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{53}$ $=(\hat{n}(j_{\beta})+\hat{n}(a_{\beta})-2\hat{n}(j_{\beta})\hat{n}(a_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{54}$ $=(\hat{n}(i_{\beta})+\hat{n}(b_{\beta})-2\hat{n}(i_{\beta})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
@@
$A_{55}$ $=(1-\hat{n}(i_{\alpha})-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(j_{\beta})+\hat{n}(i_{\alpha})\hat{n}(i_{\beta})+\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})+\hat{n}(i_{\beta})\hat{n}(j_{\beta})$
@@
$A_{56}$ $=(1-\hat{n}(a_{\alpha})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(a_{\alpha})\hat{n}(a_{\beta})+\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})+\hat{n}(a_{\beta})\hat{n}(b_{\beta})$
@@
$A_{57}$ $=(\hat{n}(j_{\beta})\hat{n}(b_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(b_{\beta})-\hat{n}(j_{\alpha})\hat{n}(j_{\beta})-\hat{n}(j_{\alpha})\hat{n}(b_{\beta})$
@@
$A_{58}$ $=(\hat{n}(i_{\alpha})\hat{n}(i_{\beta})-\hat{n}(i_{\alpha})\hat{n}(a_{\beta})+\hat{n}(i_{\beta})\hat{n}(a_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(a_{\beta})$
@@
$A_{59}$ $=-(1-\hat{n}(a_{\alpha})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(a_{\alpha})\hat{n}(a_{\beta})+\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})+\hat{n}(a_{\beta})\hat{n}(b_{\beta})$
@@
$A_{60}$ $=-(1-\hat{n}(i_{\alpha})-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(j_{\beta})+\hat{n}(i_{\alpha})\hat{n}(i_{\beta})+\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})+\hat{n}(i_{\beta})\hat{n}(j_{\beta})$
@@
$A_{61}$ $=(\hat{n}(a_{\alpha})\hat{n}(a_{\beta})-\hat{n}(i_{\alpha})\hat{n}(i_{\beta})-\hat{n}(i_{\alpha})\hat{n}(a_{\beta})+\hat{n}(i_{\beta})\hat{n}(a_{\alpha})$
@@
$A_{62}$ $=(\hat{n}(j_{\alpha})\hat{n}(j_{\beta})-\hat{n}(j_{\alpha})\hat{n}(b_{\beta})+\hat{n}(j_{\beta})\hat{n}(b_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(b_{\beta})$

### v2

@@ I Introduction
Conventional unitary coupled cluster (UCC) ansätze[1] and their adaptive extensions[13, 11, 5, 14, 3] construct the variational wavefunction from a pool of anti-Hermitian generators chosen to balance expressiveness and circuit depth. In principle, the complete set of $4^{N}-1$ traceless Pauli generators acting on $N$ qubits spans the full $\mathfrak{su}(2^{N})$ Lie algebra, guaranteeing convergence to the exact ground state within a fixed orbital basis. The exponential scaling of this set, however, makes its direct use computationally infeasible. To achieve polynomial scaling, smaller operator pools that are closed under commutation have been proposed in both fermionic[5] and qubit operator spaces[14]. While such reduced pools can in principle generate the full algebra, many of them break key physical symmetries, including total spin ($S^{2}$), its projection ($S_{z}$), and particle number ($N_{e}$).
Maintaining these symmetries during variational optimization is desirable not only for physical consistency but also for improved convergence and reduced redundancy in the variational parameters. Several strategies have been proposed to enforce or restore symmetry, including penalty-based constraints[12], projection methods[17], and modified symmetry-conserving generators[2, 3]. The DISCO framework[3], for example, introduces one-electron and electron-pair rotations that preserve total spin by construction. Although effective for spin conservation, this approach restricts the operator pool to zero-seniority excitations (pair rotations), thereby reducing the overall expressivity of the ansatz.
@@
Since the number of basis elements in the relevant Lie algebras is modest (fewer than 100), the entire factorization can be carried out efficiently in the adjoint representation using low-dimensional matrix algebra. Instead of employing symbolic Baker–Campbell–Hausdorff expansions, we reduce the reparametrization problem to a small nonlinear optimization involving matrix exponentials in the adjoint representation. This numerical treatment achieves high precision while avoiding symbolic complexity.
@@ II.1 Lie algebra framework
where $A_{k}\in\mathcal{A}$ form a Lie algebra generated by the commutator closure of the $\{A_{k}\}$ operators. The dimensionality of these algebras, $m\lesssim 84$, depends on the seniority and the specific index pattern. Each separate exponential $\exp(tA_{k})$ admits a well-known gate representation, and therefore our objective is to decompose
@@
The first question is whether such a product representation exists for the Lie algebras considered. This issue is nontrivial because certain Lie algebras contain generator sets for which the exponential of a linear combination cannot be represented as a finite product of exponentials of the individual generators [4]. In our case, all generators are anti-Hermitian operators that form subalgebras of the compact $\mathfrak{su}(N)$ algebra, for which the exponential map is surjective. Consequently, any element of the Lie group generated by the left-hand side (LHS) of Eq. (1) can be expressed by an appropriate choice of parameters in the right-hand side (RHS).
@@ II.3 Adjoint representation
To determine the parameters in Eq. (1) we use the adjoint representation, which maps each algebra element to a matrix of dimension equal to that of the Lie algebra. For the algebras considered here, the dimensions (depending on seniority and index pattern) are $5$, $28$, and $84$. These sizes are substantially smaller than those arising in faithful representations built in the computational basis of the corresponding spin orbitals. For example, using the computational basis in the Fock space of the involved spin orbitals requires matrices of dimensions $64$ and $256$ for seniorities $2$ and $4$, respectively. Therefore, we adopt the adjoint representation, in which operator matrices are constructed from the commutator (structure) constants: each element $A_{i}$ corresponds to a matrix $M^{(i)}$ defined by
@@
This objective is representation-invariant within $\exp(\mathfrak{s})$ and encodes the Baker–Campbell–Hausdorff identity for the chosen semisimple Lie algebra. In practice, we employ a smooth quasi-Newton optimizer (BFGS). Because $m\leq 24$ in all cases considered, each evaluation of the norm difference is inexpensive and convergence is rapid. Code to perform this minimisation is available on GitHub[8].
@@ II.4 Implementation of unitaries
$N_{\text{Pauli}}^{\text{JW}}$ 48 48 384 640
@@
To reduce the Pauli-string count, the generators ($A_{s}$) are grouped into groups of commuting operators. For all of the three cases, seniority-2, seniority-4 (intermediate singlet), and, seniority-4 (intermediate triplet), this results in three groups. The groups of commuting generators for the seniority-2 case are,
$s$ $=\{1,5\},\quad s=\{2,4\},\quad s=\{3\}$ (19)
The groups for the seniority-4 cases can be found in Appendices AB.
@@ III.1 Computational details
To evaluate the effect of spin-adapted fermionic generators on the efficiency and accuracy of variational algorithms, we carried out a series of adaptive VQE calculations using different operator pools. All calculations employed the STO-3G[7, 6] basis set and were converged when the maximum absolute value of the gradient of the operator pool was below $10^{-5}$ a.u. The state-vector simulator implemented in SlowQuant[9] was used to perform noiseless classical emulations. Energy minimizations employed the BFGS algorithm as implemented in SciPy[16].
The operator pool for the conventional fermionic singles and doubles (SD) adaptive fermionic ansatz consisted of $\{\hat{G}_{i_{\sigma}}^{a_{\sigma}},\,\hat{G}_{i_{\sigma}j_{\tau}}^{a_{\sigma}b_{\tau}},\,\hat{G}_{i_{\sigma}j_{\tau}}^{a_{\tau}b_{\sigma}}\},\quad i,j<a,b$, this ansatz is equal to the ADAPT ansatz[5]. For the spin-adapted fermionic ansatz ($\mathrm{S}_{\mathrm{SA}}\mathrm{D}_{\mathrm{SA}}$), the operator pool was $\{\hat{G}^{\mathrm{SA}}_{ia},\,\hat{G}^{\mathrm{SA}}_{ijab},\,\hat{G}^{\mathrm{SA}^{\prime}}_{ijab}\},\quad i,j<a,b$. Here, $\hat{G}^{\mathrm{SA}}_{ia}=\hat{G}_{i_{\alpha}}^{a_{\alpha}}+\hat{G}_{i_{\beta}}^{a_{\beta}}$, is a fermionic spin-adapted single excitation operator. Finally, for the spin-adapted singles and pair-doubles ansatz ($\mathrm{S}_{\mathrm{SA}}\mathrm{D}_{\mathrm{pair}}$), the operator pool was $\{\hat{G}^{\mathrm{SA}}_{ia},\,\hat{G}^{\mathrm{SA}}_{iiaa}\},\quad i<a,$ where $\hat{G}^{\mathrm{SA}}_{iiaa}$ corresponds to $\hat{G}_{i_{\alpha}i_{\beta}}^{a_{\alpha}a_{\beta}}$. Each adaptive VQE ansatz was constructed iteratively by selecting the operator with the largest energy gradient magnitude at each step until the convergence criterion was reached.
@@ III.2 Adaptive VQE calculations
Figure 1: Adaptive VQE calculations for $\text{H}_{2}\text{O}$ with $\theta_{\mathrm{HOH}}=109^{\circ}$ and $R_{\mathrm{OH}}=0.977$ Å. SD denotes the fermionic operator pool, $\text{S}_{\text{SA}}\text{D}_{\text{SA}}$ the spin-adapted fermionic operator pool, and $\text{S}_{\text{SA}}\text{D}_{\text{pair}}$ the pool of spin-adapted singles and pair doubles.
@@
The alternative pool of spin-adapted singles and pair doubles (green dashed line) also conserves total spin and consists of circuits that can be directly implemented with known gate structures. However, this ansatz frequently becomes trapped in local minima and, in this case, converged to a solution $8.8\times 10^{-3}$ Hartree above the FCI reference. A similar shortcoming of pair-doubles-based operator pools has been reported previously by Burton et al.[3]
Figure 2: Adaptive VQE calculations for $\text{Be}\text{H}_{2}$ with $\theta_{\mathrm{HBeH}}=180^{\circ}$ and $R_{\mathrm{BeH}}=1.35$ Å. SD denotes the fermionic operator pool, $\text{S}_{\text{SA}}\text{D}_{\text{SA}}$ the spin-adapted fermionic operator pool, and $\text{S}_{\text{SA}}\text{D}_{\text{pair}}$ the pool of spin-adapted singles and pair doubles.
@@
Previous studies have attempted to address this problem by adding penalty terms to the Hamiltonian during variational optimization[10] or by applying spin-projection techniques[15]. However, these approaches increase computational cost either by complicating the energy functional or by requiring additional ancilla qubits. Using spin-adapted operators naturally preserves the desired spin symmetry without modifying the Hamiltonian or introducing extra qubits, offering a simple and physically consistent solution.
@@ Appendix B Seniority 4, intermediate triplet
This 84-dimensional Lie algebra originates from the commutator closure of 6 elementary generators (the closure produces 84 linearly independent as fermionic operators)
@@
$A_{41}$ $=(\hat{n}(j_{\alpha})+\hat{n}(b_{\alpha})-2\hat{n}(j_{\alpha})\hat{n}(b_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{42}$ $=(\hat{n}(i_{\alpha})\hat{n}(j_{\beta})-\hat{n}(i_{\alpha})\hat{n}(b_{\beta})-\hat{n}(j_{\beta})\hat{n}(a_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{43}$ $=(\hat{n}(i_{\alpha})+\hat{n}(a_{\alpha})-2\hat{n}(i_{\alpha})\hat{n}(a_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\beta}b_{\alpha}}$
$A_{44}$ $=(\hat{n}(i_{\beta})\hat{n}(j_{\alpha})-\hat{n}(i_{\beta})\hat{n}(b_{\alpha})-\hat{n}(j_{\alpha})\hat{n}(a_{\beta})+\hat{n}(a_{\beta})\hat{n}(b_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\alpha}b_{\beta}}$
$A_{45}$ $=(\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})-2\hat{n}(i_{\alpha})\hat{n}(b_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
$A_{46}$ $=(\hat{n}(i_{\beta})\hat{n}(j_{\alpha})-\hat{n}(i_{\beta})\hat{n}(a_{\alpha})-\hat{n}(j_{\alpha})\hat{n}(b_{\beta})+\hat{n}(a_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{47}$ $=(\hat{n}(j_{\alpha})+\hat{n}(a_{\alpha})-2\hat{n}(j_{\alpha})\hat{n}(a_{\alpha}))\hat{G}_{i_{\alpha}j_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{48}$ $=(\hat{n}(i_{\alpha})\hat{n}(j_{\beta})-\hat{n}(i_{\alpha})\hat{n}(a_{\beta})-\hat{n}(j_{\beta})\hat{n}(b_{\alpha})+\hat{n}(a_{\beta})\hat{n}(b_{\alpha}))\hat{G}_{i_{\beta}j_{\alpha}}^{a_{\alpha}b_{\beta}}$
@@
$A_{49}$ $=(1-\hat{n}(i_{\alpha})-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(j_{\beta})+\hat{n}(i_{\alpha})\hat{n}(i_{\beta})+\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})+\hat{n}(i_{\beta})\hat{n}(j_{\beta})$
@@
$A_{50}$ $=(1-\hat{n}(a_{\alpha})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(a_{\alpha})\hat{n}(a_{\beta})+\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})+\hat{n}(a_{\beta})\hat{n}(b_{\beta})$
@@
$A_{51}$ $=(\hat{n}(j_{\beta})\hat{n}(b_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(b_{\beta})-\hat{n}(j_{\alpha})\hat{n}(j_{\beta})-\hat{n}(j_{\alpha})\hat{n}(b_{\beta})$
@@
$A_{52}$ $=(\hat{n}(i_{\alpha})\hat{n}(i_{\beta})-\hat{n}(i_{\alpha})\hat{n}(a_{\beta})+\hat{n}(i_{\beta})\hat{n}(a_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(a_{\beta})$
@@
$A_{53}$ $=-(1-\hat{n}(a_{\alpha})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(a_{\alpha})\hat{n}(a_{\beta})+\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})+\hat{n}(a_{\beta})\hat{n}(b_{\beta})$
@@
$A_{54}$ $=-(1-\hat{n}(i_{\alpha})-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(j_{\beta})+\hat{n}(i_{\alpha})\hat{n}(i_{\beta})+\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})+\hat{n}(i_{\beta})\hat{n}(j_{\beta})$
@@
$A_{55}$ $=(\hat{n}(a_{\alpha})\hat{n}(a_{\beta})-\hat{n}(i_{\alpha})\hat{n}(i_{\beta})-\hat{n}(i_{\alpha})\hat{n}(a_{\beta})+\hat{n}(i_{\beta})\hat{n}(a_{\alpha})$
@@
$A_{56}$ $=(\hat{n}(j_{\alpha})\hat{n}(j_{\beta})-\hat{n}(j_{\alpha})\hat{n}(b_{\beta})+\hat{n}(j_{\beta})\hat{n}(b_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(b_{\beta})$